OpenStudy (anonymous):

Assuming the height of a redwood to be 105m ( about 350 ft), estimate the osmotic pressure required to push water up to the tree top?

OpenStudy (aaronq):

\[\Pi = pgh\]

OpenStudy (anonymous):

what's p

OpenStudy (aaronq):

density

OpenStudy (aaronq):

g is gravity

OpenStudy (anonymous):

i did that but still got it wrong

mathslover (mathslover):

@cleverchuck , welcome to openstudy. Please let us know what you tried yet in solving the question, the formula given by @aaronq should work, what values did you put on the place of 'p' , 'g' and 'h' ?

mathslover (mathslover):

Well first of all , what is osmotic pressure? Osmotic Pressure \((\bf \pi)\) is the pressure required to prevent the osmosis from occurring. \(\bf \pi\) = \(\rho g h\) \(\rho = \textbf{density}\) \(g = \textbf{ acceleration due to gravity} \) \(h = \textbf{height (here it is given)}\) Now, you know the density of the water , (but note , you have to put the value according to the units as the height is given in metres , so you have to put the value of the density of water having unit \(kg m^{-3} \) ) the acceleration due to gravity = g (I hope you know the value of it) and the "h" that is given here. (105 metres )

mathslover (mathslover):

Just solve for it and it must be right.